OpenStudy (anonymous):

what is the sipmlest form of 3/11 times 5/12

OpenStudy (anonymous):

\[\frac{3}{11}\times\frac{5}{12}=\frac{3}{11}\times\frac{5}{3(4)}=\frac{1}{11}\times\frac{5}{4}=\frac{5}{44}\]

OpenStudy (anonymous):

15 / 132 Both divisible by 3 = 5 / 44

OpenStudy (anonymous):

That is a useful method jcari. I just like doing it the top way because I can automatically see that I can factor out a \(3\)

OpenStudy (anonymous):

ohh I see what you did now, I was trying to figure it out. Good job, need to use that : )

OpenStudy (anonymous):

Haha

OpenStudy (anonymous):

Thank you